# Predicting perturbation targets with causal differential networks

### v1

@@ B.1 Biological datasets
Figure 5: Heatmap of correlation between log-fold change, sorted by cluster.
Figure 6: Perturb-seq dataset statistics, after processing (Table 6).
Figure 7: Sci-Plex dataset statistics, after processing. There are 3 chemical perturbations per cell line.
@@ B.2 Synthetic data
Root causal mechanisms are uniform. For hard interventions, we set $f(x)\leftarrow z$, where $z\sim\text{Uniform}(-1,1)$. For soft interventions, we set
@@
Appendix C Implementation details
C.1 Baselines
We used the latest releases of all baselines. For GenePt, this corresponds to the v2 March 2024 update, which used newer models and additional protein data, compared to their initial paper (GenePT_gene_protein_embedding_model_3_text).
Due to the varying coverage of genes in their domain knowledge, several baselines failed to make predictions for certain genes. Of the 2,842 unique genetic perturbations, only 2,819 ($99.1\%$) mapped to GenePt embeddings. The remainder used the mean gene embedding (within the dataset) as the language-based embedding.
Figure 8 depicts the low, but non-trivial proportion of differentially expressed genes for which Gears was unable to make a prediction, due to lack of node coverage in their processed gene ontology graph (Ashburner et al., 2000). These genes were not considered in the rankings or evaluations for Gears.
Appendix D Additional analyses
Table 7: Ablation studies. We re-train Cdn without pretrained aggregator weights, marginal estimates, or global statistics.
K562 gw K562 es RPE1 HepG2 Jurkat
Model rank $R_{20}$ $r$ rank $R_{20}$ $r$ rank $R_{20}$ $r$ rank $R_{20}$ $r$ rank $R_{20}$ $r$
Cdn 0.72 0.37 0.32 0.56 0.15 0.35 0.59 0.12 0.53 0.55 0.07 0.40 0.62 0.10 0.39
--pretrain $0.50$ $0.14$ $0.20$ $0.51$ $0.01$ $0.23$ $0.49$ $0.03$ $0.50$ $0.52$ $0.03$ $0.37$ $0.54$ $0.07$ $0.30$
--marginal $0.55$ $0.15$ $0.19$ $0.49$ $0.06$ $0.25$ $0.50$ $0.05$ $0.47$ $0.49$ $0.05$ $0.39$ $0.58$ $0.04$ $0.31$
--global $0.52$ $0.14$ $0.20$ $0.49$ $0.04$ $0.27$ $0.47$ $0.04$ $0.49$ $0.49$ $0.04$ $0.35$ $0.46$ $0.06$ $0.25$
Table 8: Ablation study on global statistics over synthetic datasets. Mlp and Axial use inverse covariance (same as pretraining). Corr was finetuned for correlation, and Corr+Cov was finetuned for correlation and covariance (concatenated).
$N$ $E$ Cdn Variant Linear (Hard) Linear (Soft) Poly. (Hard) Poly. (Soft)
[…]
Axial $\textbf{0.73}\scriptstyle\pm.05$ $\textbf{0.85}\scriptstyle\pm.03$ $\textbf{0.48}\scriptstyle\pm.06$ $\textbf{0.64}\scriptstyle\pm.05$ $\textbf{0.62}\scriptstyle\pm.09$ 0.77±.07plus-or-minus0.77.070.$7$07 $\textbf{0.73}\scriptstyle\pm.07$ $\textbf{0.82}\scriptstyle\pm.06$
+Corr $0.71\scriptstyle\pm.07$ $0.84\scriptstyle\pm.04$ $0.32\scriptstyle\pm.05$ $0.47\scriptstyle\pm.07$ $0.58\scriptstyle\pm.08$ $0.74\scriptstyle\pm.05$ $0.27\scriptstyle\pm.04$ $0.34\scriptstyle\pm.04$
+Corr+Cov $0.68\scriptstyle\pm.04$ $0.82\scriptstyle\pm.02$ $0.39\scriptstyle\pm.03$ $0.50\scriptstyle\pm.03$ $0.58\scriptstyle\pm.07$ $0.71\scriptstyle\pm.07$ $0.48\scriptstyle\pm.05$ $0.65\scriptstyle\pm.05$
[…]
+Corr $\textbf{0.78}\scriptstyle\pm.03$ 0.88±.02plus-or-minus0.88.020.$8$02 $0.35\scriptstyle\pm.02$ $0.52\scriptstyle\pm.02$ $0.73\scriptstyle\pm.06$ $0.84\scriptstyle\pm.03$ $0.34\scriptstyle\pm.05$ $0.41\scriptstyle\pm.06$
+Corr+Cov $0.75\scriptstyle\pm.03$ $0.84\scriptstyle\pm.02$ $0.34\scriptstyle\pm.03$ $0.47\scriptstyle\pm.06$ $0.69\scriptstyle\pm.07$ $0.79\scriptstyle\pm.05$ $0.53\scriptstyle\pm.04$ $\textbf{0.68}\scriptstyle\pm.04$
20 20 Mlp $\textbf{0.67}\scriptstyle\pm.05$ $\textbf{0.87}\scriptstyle\pm.02$ $0.28\scriptstyle\pm.01$ $0.54\scriptstyle\pm.03$ $60$06 $\textbf{0.84}\scriptstyle\pm.03$ $0.34\scriptstyle\pm.04$ $0.51\scriptstyle\pm.04$
Axial $0.65\scriptstyle\pm.03$ $0.86\scriptstyle\pm.01$ $\textbf{0.49}\scriptstyle\pm.05$ $\textbf{0.72}\scriptstyle\pm.03$ $0.58\scriptstyle\pm.05$ $0.83\scriptstyle\pm.02$ $\textbf{0.65}\scriptstyle\pm.02$ $\textbf{0.81}\scriptstyle\pm.02$
+Corr $0.62\scriptstyle\pm.03$ $0.83\scriptstyle\pm.01$ $0.25\scriptstyle\pm.05$ $0.47\scriptstyle\pm.05$ $0.58\scriptstyle\pm.05$ $0.81\scriptstyle\pm.02$ $0.21\scriptstyle\pm.02$ $0.37\scriptstyle\pm.03$
+Corr+Cov $0.60\scriptstyle\pm.03$ $0.81\scriptstyle\pm.01$ $0.31\scriptstyle\pm.03$ $0.58\scriptstyle\pm.01$ $0.58\scriptstyle\pm.04$ $0.80\scriptstyle\pm.02$ $0.40\scriptstyle\pm.03$ $0.65\scriptstyle\pm.03$
[…]
Axial $0.74\scriptstyle\pm.05$ $0.90\scriptstyle\pm.02$ $\textbf{0.38}\scriptstyle\pm.03$ $\textbf{0.65}\scriptstyle\pm.01$ $0.68\scriptstyle\pm.08$ $0.87\scriptstyle\pm.04$ $\textbf{0.64}\scriptstyle\pm.03$ $\textbf{0.79}\scriptstyle\pm.03$
[…]
+Corr+Cov $0.73\scriptstyle\pm.05$ $0.87\scriptstyle\pm.02$ $0.36\scriptstyle\pm.03$ $0.58\scriptstyle\pm.03$ $0.62\scriptstyle\pm.07$ $0.82\scriptstyle\pm.03$ $0.43\scriptstyle\pm.04$ $0.67\scriptstyle\pm.02$
Ablation studies
We run ablation studies on the model architecture (Table 7) by re-training Cdn without pretrained aggregator weights, global statistics, or marginal estimates. We find that pretrained weights and global correlation are essential to model performance. The marginal estimates are of mixed importance: they are beneficial on the genome-wide (gw) dataset and Jurkat, but less on the others.
Table 8 ablates variants of Cdn trained with different global statistics (but initialized with the inverse covariance pretrained weights) on synthetic data. The default summary statistic (Mlp, Axial) is inverse covariance. Corr denotes the correlation version, later finetuned on Perturb-seq data (since correlation is much easier to compute on large graphs, compared to inverse covariance). The Corr+Cov version aligns exactly with our theoretical intuition, and performs well in most settings (polynomial soft is the exception). This model modified the global layer to take an input of size $N\times N\times 2$, where each embedding was initialized to a distinct copy of the original. However, the inverse covariance version still performed the best, perhaps since there is no mismatch between the pretraining and finetuning statistics.
@@ Appendix D Additional analyses
We also compared the runtimes of various algorithms on the Perturb-seq and synthetic datasets. On the Perturb-seq data, all models finished running within minutes with the exception of Gears (Figure 9). Runtimes of the various causal algorithms varied significantly (Table 9). The slowest method was Dcdi, with an average runtime of around 10 hours on $N=20$ datasets, while the fastest were Ut-Igsp and the Mlp variant of Cdn. All models were benchmarked on equivalent hardware (A6000 GPU, 1 CPU core).
@@
Due to space limitations, we report uncertainty estimates in Table 10. Multiple baselines produce deterministic results (Linear, GenePt), so instead of model randomness, we report uncertainties over the sampling of single cells. Specifically, for each perturbation with $M$ cells, we sample
@@
Table 11 reports results on all synthetic test datasets, averaging over all interventions for each dataset.
@@ 20 Ut-Igsp 1 78 19 26
Table 10: Uncertainty quantification on Perturb-seq datasets by sub-sampling to $80\%$ of cells per perturbation (or 50, whichever is higher). Standard deviation reported over 5 rounds of sub-sampling. Cdn is also the most robust to sub-sampling, compared to baselines. GenePt performance is highly variable.
@@
Model rank $R_{20}$ $r$ rank $R_{20}$ $r$ rank $R_{20}$ $r$
Linear $0.50\scriptstyle\pm.00$ $0.17\scriptstyle\pm.01$ $0.20\scriptstyle\pm.01$ $0.50\scriptstyle\pm.01$ $10$01 $0.28\scriptstyle\pm.01$ $0.48\scriptstyle\pm.00$ $0.07\scriptstyle\pm.01$ $0.49\scriptstyle\pm.01$
Mlp $0.48\scriptstyle\pm.00$ $0.21\scriptstyle\pm.00$ $0.20\scriptstyle\pm.00$ $0.48\scriptstyle\pm.00$ 0.11±.00plus-or-minus0.11.000.$1$00 $0.19\scriptstyle\pm.00$ $0.53\scriptstyle\pm.00$ $0.15\scriptstyle\pm.00$ $0.39\scriptstyle\pm.00$
GenePt $0.54\scriptstyle\pm.09$ $0.28\scriptstyle\pm.06$ $0.29\scriptstyle\pm.08$ $0.50\scriptstyle\pm.08$ $0.21\scriptstyle\pm.09$ $0.39\scriptstyle\pm.10$ $0.51\scriptstyle\pm.10$ $0.20\scriptstyle\pm.08$ $0.51\scriptstyle\pm.07$
Gears $0.50\scriptstyle\pm.01$ $0.13\scriptstyle\pm.01$ 0.22±.01plus-or-minus0.22.010.$2$01 $0.51\scriptstyle\pm.00$ $0.07\scriptstyle\pm.00$ $0.23\scriptstyle\pm.00$ $0.46\scriptstyle\pm.00$ $0.07\scriptstyle\pm.01$ $0.41\scriptstyle\pm.02$
PdG $0.49\scriptstyle\pm.01$ $0.21\scriptstyle\pm.01$ $0.25\scriptstyle\pm.00$ $0.50\scriptstyle\pm.01$ $0.16\scriptstyle\pm.01$ $0.32\scriptstyle\pm.01$ $0.49\scriptstyle\pm.01$ $0.08\scriptstyle\pm.01$ $0.47\scriptstyle\pm.01$
@@
HepG2 Jurkat
rank $R_{20}$ $r$ rank $R_{20}$ $r$
Linear $0.49\scriptstyle\pm.00$ $0.06\scriptstyle\pm.01$ $0.42\scriptstyle\pm.01$ $0.50\scriptstyle\pm.00$ $0.08\scriptstyle\pm.00$ $0.27\scriptstyle\pm.01$
Mlp $0.50\scriptstyle\pm.00$ $0.09\scriptstyle\pm.00$ $0.41\scriptstyle\pm.00$ $0.47\scriptstyle\pm.00$ $0.12\scriptstyle\pm.00$ $0.31\scriptstyle\pm.00$
GenePt $0.50\scriptstyle\pm.13$ $\textbf{0.19}\scriptstyle\pm.08$ $0.45\scriptstyle\pm.06$ $0.51\scriptstyle\pm.07$ 0.22±.09plus-or-minus0.22.090.$2$09 $0.40\scriptstyle\pm.11$
Gears $0.51\scriptstyle\pm.00$ 0.11±.00plus-or-minus0.11.000.$1$00 $0.45\scriptstyle\pm.00$ $0.52\scriptstyle\pm.01$ $0.08\scriptstyle\pm.00$ $0.30\scriptstyle\pm.00$
PdG $0.49\scriptstyle\pm.00$ $0.05\scriptstyle\pm.01$ $0.41\scriptstyle\pm.01$ $0.49\scriptstyle\pm.01$ 0.11±.01plus-or-minus0.11.010.$1$01 $0.34\scriptstyle\pm.00$
Cdn $\textbf{0.65}\scriptstyle\pm.01$ $0.17\scriptstyle\pm.01$ $\textbf{0.46}\scriptstyle\pm.02$ $\textbf{0.65}\scriptstyle\pm.01$ $\textbf{0.24}\scriptstyle\pm.02$ $\textbf{0.43}\scriptstyle\pm.01$
Table 11: Intervention target prediction results on synthetic datasets, extended results. Uncertainty is standard deviation over 5 distinct datasets. Metrics are averaged over all $3N$ perturbations for a given dataset (1-3 targets).
$N$ $E$ Model Linear (Hard) Linear (Soft) Poly. (Hard) Poly. (Soft)

### v2

@@ D.1 Biological datasets
Figure 5: Heatmap of correlation between log-fold change, sorted by cluster, used for data splits.
Figure 6: Difference between observational and interventional correlation matrices on K562 genome-wide. In some cases, the changes are minimal and specific; in others, they are more diffuse.
Figure 7: Difference between observational and interventional correlation matrices on Sci-Plex datasets. Effects are much more diffuse than in genetic perturbations (Figure 6).
Figure 8: Transcriptomics dataset statistics, after processing. See Table 5 for details.
@@ D.2 Synthetic data
  • Neural network, non-additive $f(x)=\text{Tanh}((X,E)W_{\text{in}})W_{\text{out}}$
  • Neural network, additive: $f(x)=\text{Tanh}(XW_{\text{in}})W_{\text{out}}+E$
@@
  • Sigmoid: $f(x)=\sum_{i=1}^{d}W_{i}\cdot\text{sigmoid}(X_{i})+E$
Linear and the neural network variants were used for training. Polynomial and sigmoid were only used for testing.
Root causal mechanisms are uniform. For hard interventions, we set $f(x)\leftarrow z$, where $z\sim\text{Uniform}(-1,1)$. For soft “scale” interventions, we set
@@
For soft “shift” interventions, we set
$f(x)$ $\leftarrow f(x)+\text{Sign}(z)\cdot z_{1}$
$z_{1}$ $\sim\text{Uniform}(2,4)$
$z$ $\sim\text{Uniform}(-1,1).$
That is, we translate $f(x)$ by a scaling factor that is equal probability $\lessgtr 0$ (constant across all observations).
Appendix E Additional analyses
Figure 10: Ablation study: We compare Cdn trained on synthetic data, to the same model without any ground truth graph information (no pretraining or $\mathcal{L}_{G}$). We find that graph information leads to more stable and better results. Validation metrics and loss. One step is one batch (16 samples).
@@ Appendix E Additional analyses
We compared the runtimes of various algorithms on the Perturb-seq and synthetic datasets. On the Perturb-seq data, all models finished running within minutes with the exception of Gears (Figure 11). Runtimes of the various causal algorithms varied significantly (Table 11). The slowest method was Dcdi, with an average runtime of around 10 hours on $N=20$ datasets, while the fastest were Ut-Igsp and the Mlp variant of Cdn. All models were benchmarked on equivalent hardware (A6000 GPU, 1 CPU core).
@@
Due to space limitations, we report uncertainty estimates in Table 7. Multiple baselines produce deterministic results (Linear, Dge, GenePt), so instead of model randomness, we report uncertainties over the sampling of single cells. Specifically, for each perturbation with $M$ cells, we sample
@@
Perturb-seq trivial perturbations
For the main results in Table 1, we focus on genetic perturbations for which the intended gene target was not the gene with the greatest log-fold change. This is because genetic perturbations are (unrealistically) specific, compared to drugs or transcription factors. In Table 8, we show that our model performs nearly perfectly in this easy setting.
Graph loss
We assess the value of predicting the causal graph, as opposed to only training a supervised classifier (Figure 10). Without modifying the architecture (which has inductive biases towards the graph modality), we train a version of Cdn without any graph labels (causal graph learner no longer pretrained; remove graph loss $\mathcal{L}_{G}$). Across the validation loss and various metrics, we find that the full Cdn is more stable, and leads to better results. This model corresponds to the ”no $\mathcal{L}_{G}$” test results in Table 3.
@@
During model development, we realized that if the model is allowed to train on all types of perturbations, it can easily overfit and learn these distributional shifts very well. For example, in Table 9, the models trained on both types of soft interventions performs perfectly (unrealistic) in many cases, even though the graphs and causal mechanism parameters are completely novel. Therefore, to assess whether our model can generalize, we chose to hold out a class of soft intervention for our main results in Table 3.
Table 10 reports results on all synthetic test datasets, averaging over all interventions for each dataset. We also include the “full” version of Cdn, i.e. trained on all mechanisms, in this analysis.
@@ 20 Ut-Igsp 1 78 19 26
Table 7: Uncertainty quantification on Perturb-seq datasets, seen cell lines, by sub-sampling to $80\%$ of cells per perturbation (or 50, whichever is higher). Note that results may be slightly different from Table 1 due to sub-sampling. Standard deviation reported over 5 seeds. GenePt performance is highly variable. Dge is quite sensitive to sub-sampling on K562 genome-wide.
@@
Model rank $\rho$ $r$ rank $\rho$ $r$ rank $\rho$ $r$
Linear $0.50\scriptstyle\pm.00$ $0.17\scriptstyle\pm.01$ $0.20\scriptstyle\pm.01$ $0.50\scriptstyle\pm.01$ $0.25\scriptstyle\pm.01$ $0.28\scriptstyle\pm.01$ $0.48\scriptstyle\pm.00$ $0.40\scriptstyle\pm.01$ $0.49\scriptstyle\pm.01$
Mlp $0.48\scriptstyle\pm.00$ $0.17\scriptstyle\pm.00$ $0.20\scriptstyle\pm.00$ $0.48\scriptstyle\pm.00$ $0.16\scriptstyle\pm.00$ $0.19\scriptstyle\pm.00$ $0.53\scriptstyle\pm.00$ $0.30\scriptstyle\pm.00$ $0.39\scriptstyle\pm.00$
Dge $0.72\scriptstyle\pm.01$ $0.48\scriptstyle\pm.02$ $0.52\scriptstyle\pm.02$ $0.86\scriptstyle\pm.00$ $0.59\scriptstyle\pm.02$ $0.61\scriptstyle\pm.02$ $0.74\scriptstyle\pm.00$ $0.56\scriptstyle\pm.00$ $0.64\scriptstyle\pm.00$
GenePt $0.54\scriptstyle\pm.09$ $0.26\scriptstyle\pm.08$ $0.29\scriptstyle\pm.08$ $0.50\scriptstyle\pm.08$ $0.36\scriptstyle\pm.10$ $0.39\scriptstyle\pm.10$ $0.51\scriptstyle\pm.10$ $0.43\scriptstyle\pm.08$ $0.51\scriptstyle\pm.07$
Gears $0.50\scriptstyle\pm.01$ $0.19\scriptstyle\pm.01$ 0.22±.01plus-or-minus0.22.010.$2$01 $0.51\scriptstyle\pm.00$ $0.20\scriptstyle\pm.01$ $0.23\scriptstyle\pm.00$ $0.46\scriptstyle\pm.00$ 0.33±.01plus-or-minus0.33.010.$3$01 $0.41\scriptstyle\pm.02$
PdG $0.49\scriptstyle\pm.01$ 0.22±.00plus-or-minus0.22.000.$2$00 $0.25\scriptstyle\pm.00$ $0.50\scriptstyle\pm.01$ $0.28\scriptstyle\pm.01$ $0.32\scriptstyle\pm.01$ $0.49\scriptstyle\pm.01$ $0.38\scriptstyle\pm.01$ $0.47\scriptstyle\pm.01$
Cdn $\textbf{0.91}\scriptstyle\pm.01$ $\textbf{0.71}\scriptstyle\pm.02$ $\textbf{0.72}\scriptstyle\pm.02$ $\textbf{0.95}\scriptstyle\pm.00$ $\textbf{0.80}\scriptstyle\pm.03$ $\textbf{0.81}\scriptstyle\pm.03$ $\textbf{0.92}\scriptstyle\pm.00$ $\textbf{0.65}\scriptstyle\pm.02$ $\textbf{0.74}\scriptstyle\pm.01$
HepG2 Jurkat
rank $\rho$ $r$ rank $\rho$ $r$
Linear $0.49\scriptstyle\pm.00$ $0.34\scriptstyle\pm.01$ $0.42\scriptstyle\pm.01$ $0.50\scriptstyle\pm.00$ $0.23\scriptstyle\pm.01$ $0.27\scriptstyle\pm.01$
Mlp $0.50\scriptstyle\pm.00$ $0.34\scriptstyle\pm.00$ $0.41\scriptstyle\pm.00$ $0.47\scriptstyle\pm.00$ $0.25\scriptstyle\pm.00$ $0.31\scriptstyle\pm.00$
Dge $0.83\scriptstyle\pm.00$ $0.49\scriptstyle\pm.01$ $0.50\scriptstyle\pm.01$ $0.78\scriptstyle\pm.01$ 0.44±.02plus-or-minus0.44.020.$4$02 $0.49\scriptstyle\pm.02$
GenePt $0.50\scriptstyle\pm.13$ $0.39\scriptstyle\pm.07$ $0.45\scriptstyle\pm.06$ $0.51\scriptstyle\pm.07$ $0.36\scriptstyle\pm.11$ $0.40\scriptstyle\pm.11$
Gears $0.51\scriptstyle\pm.00$ $0.38\scriptstyle\pm.00$ $0.45\scriptstyle\pm.00$ $0.52\scriptstyle\pm.01$ $0.24\scriptstyle\pm.00$ $0.30\scriptstyle\pm.00$
PdG $0.49\scriptstyle\pm.00$ $0.34\scriptstyle\pm.01$ $0.41\scriptstyle\pm.01$ $0.49\scriptstyle\pm.01$ $0.29\scriptstyle\pm.00$ $0.34\scriptstyle\pm.00$
Cdn $\textbf{0.93}\scriptstyle\pm.00$ $\textbf{0.67}\scriptstyle\pm.01$ $\textbf{0.73}\scriptstyle\pm.01$ $\textbf{0.96}\scriptstyle\pm.01$ $\textbf{0.83}\scriptstyle\pm.04$ $\textbf{0.85}\scriptstyle\pm.03$
Table 8: Results on 5 Perturb-seq datasets, trivial perturbations (intended target has largest log-fold change, though not necessarily smallest adjusted p-value). Top: Train on all cell lines jointly. Bottom: Leave one cell line out of training. Test sets (unseen perturbations) are identical in both settings. Cdn (synthetic) and Dge do not require training on real Perturb-seq data. Metrics, from left to right: normalized rank of ground truth, top 1 Spearman and Pearson correlations.
K562 (gw) K562 (es) RPE1 HepG2 Jurkat
Setting Model rank $\rho$ $r$ rank $\rho$ $r$ rank $\rho$ $r$ rank $\rho$ $r$ rank $\rho$ $r$
Seen cell line Linear $0.49$ $0.18$ $0.21$ $0.50$ $0.28$ $0.31$ $0.48$ $0.37$ $0.45$ $0.48$ $0.34$ $0.43$ $0.48$ $0.24$ $0.29$
Mlp $0.48$ $0.16$ $0.20$ $0.47$ $0.17$ $0.21$ $0.54$ $0.29$ $0.37$ $0.51$ $0.35$ $0.42$ $0.48$ $0.26$ $0.32$
GenePt $0.60$ $0.36$ $0.39$ $0.47$ $0.23$ $0.26$ $0.57$ $0.49$ $0.55$ $0.44$ $0.35$ $0.41$ $0.55$ $0.48$ $0.53$
Gears $0.49$ $0.15$ $0.18$ $0.51$ $0.25$ $0.27$ $0.49$ $0.32$ $0.40$ $0.48$ $0.32$ $0.40$ $0.47$ $0.20$ $0.24$
PdG $0.49$ $0.23$ $0.26$ $0.52$ $0.31$ $0.34$ $0.51$ $0.34$ $0.42$ $0.46$ $0.34$ $0.40$ $0.49$ $0.30$ $0.35$
Cdn (Perturb-seq) 0.99 0.97 0.97 0.99 0.98 0.98 0.99 0.93 0.94 0.99 0.95 0.96 0.99 0.96 0.96
Unseen cell line Dge $0.87$ $0.68$ $0.70$ $0.92$ $0.77$ $0.79$ $0.85$ $0.73$ $0.77$ $0.89$ $0.59$ $0.60$ $0.91$ $0.73$ $0.77$
PdG $0.40$ $0.12$ $0.15$ $0.41$ $0.22$ $0.24$ $0.48$ $0.28$ $0.37$ $0.47$ $0.33$ $0.41$ $0.53$ $0.28$ $0.33$
Cdn (synthetic) $0.91$ $0.72$ $0.74$ $0.92$ $0.83$ $0.84$ $0.94$ $0.83$ $0.86$ $0.90$ $0.73$ $0.77$ $0.95$ $0.82$ $0.84$
Cdn (Perturb-seq) 0.98 0.95 0.95 0.99 0.95 0.96 0.99 0.92 0.94 0.99 0.95 0.96 0.99 0.96 0.96
Table 9: Intervention target prediction results on synthetic datasets with $N=20,E=40$. Top: hard interventions (uniform); bottom: soft interventions (scale). We found that if the model is allowed to see the class of soft intervention (“all mechanisms”) during training, it can easily learn (overfit) the distribution shift perfectly, even though test datasets and graphs are novel. Thus, our main results in Table 3 treat “scale” as an unseen soft intervention.
Linear (1) Linear (3) Polynomial (1) Polynomial (3)
@@
Hard all mechanisms ($h_{\text{cat}}$) $0.86\scriptstyle\pm.09$ $0.97\scriptstyle\pm.02$ $0.91\scriptstyle\pm.06$ $0.97\scriptstyle\pm.03$ $0.86\scriptstyle\pm.04$ $0.97\scriptstyle\pm.01$ $0.91\scriptstyle\pm.03$ $0.97\scriptstyle\pm.01$
all mechanisms ($h_{\text{diff}}$) $0.78\scriptstyle\pm.05$ $0.95\scriptstyle\pm.03$ $0.89\scriptstyle\pm.04$ $0.96\scriptstyle\pm.02$ $0.83\scriptstyle\pm.05$ $0.96\scriptstyle\pm.01$ $0.89\scriptstyle\pm.03$ $0.96\scriptstyle\pm.02$
exclude scale $0.82\scriptstyle\pm.05$ $0.96\scriptstyle\pm.02$ 0.88±.04plus-or-minus0.88.040.$8$04 $0.95\scriptstyle\pm.02$ $0.85\scriptstyle\pm.05$ $0.97\scriptstyle\pm.01$ $0.90\scriptstyle\pm.04$ $0.96\scriptstyle\pm.02$
Soft all mechanisms ($h_{\text{cat}}$) $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$ 0.99±.01plus-or-minus0.99.010.$9$01 $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$
all mechanisms ($h_{\text{diff}}$) 0.99±.01plus-or-minus0.99.010.$9$01 $1.0\scriptstyle\pm.00$ $0.98\scriptstyle\pm.01$ 0.99±.00plus-or-minus0.99.000.$9$00 $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$
exclude scale $0.83\scriptstyle\pm.10$ $0.95\scriptstyle\pm.05$ $0.76\scriptstyle\pm.07$ $0.89\scriptstyle\pm.04$ $0.97\scriptstyle\pm.03$ $1.0\scriptstyle\pm.00$ $0.96\scriptstyle\pm.02$ 0.99±.01plus-or-minus0.99.010.$9$01
Table 10: Intervention target prediction results on synthetic datasets, extended results. Uncertainty is standard deviation over 5 i.i.d. datasets. Metrics are averaged over all $3N$ perturbations for a given dataset (1-3 targets). Cdn (full) denotes that we trained on all intervention mechanisms jointly, i.e. as in Table 9.
$N$ $E$ Model Linear (Hard) Linear (Soft) Poly. (Hard) Poly. (Soft) Sigmoid. (Hard) Sigmoid (Soft)